OpenStudy (anonymous):

a right triangle has one leg with a length of 48 and hypotenuse with a length of 80. what is the length of the leg? a. 63 b. 64 c. 65 d. 66

OpenStudy (anonymous):

64. 48^2 + b^2= 80^2 2304+b^2=6400 b^2=4096 b=64 2304+4096=6400